OpenStudy (anonymous):

In a 30°- 60° - 90° right triangle, the shorter leg has a length of 4. What is the length of the longer leg?

OpenStudy (anonymous):

us the definition of sin theta = opp/hyp thus hyp = opp/sin theta.

OpenStudy (anonymous):

\[\sqrt{3}^{2} + 4^{2} = 19\] therefore, \[\sqrt{19} = 4.358898944 \] Btw, I am not sure, if this is correct